A gas has a pressure of 100 kPa at a volume of 20 L. If the final volume is 15 L, what is the final pressure (kPa)?
1
Expert's answer
2020-05-05T03:37:26-0400

 Boyle's Law: p1*V1=p2*V2;

p2=p1*V1/V2=100*20/15=133.33 kPa

Answer: the final pressure is 133.33 kPa